OpenStudy (anonymous):

A jar of mayonnaise is shaped like a cylinder. The radius of the jar is 1.5 in. and its height is 7 in. The mayonnaise is on sale for $0.08 per in3. What is the cost of the jar of mayonnaise, rounded to the nearest cent? Use 3.14 to approximate pi. $______

OpenStudy (shamim):

Welcome to openstudy

OpenStudy (shamim):

Anyway, cross sectional area of ur cylinder A=pi*r^2=3.14*1.5^2

OpenStudy (shamim):

Volume of ur cylinder=pi*r^2*h=3.14*1.5^2*7=?in^3

OpenStudy (shamim):

Total cost=0.08*volume of cylinder

OpenStudy (shamim):

Response plz!

OpenStudy (anonymous):

sorry so 0.08? @shamim

OpenStudy (shamim):

No

OpenStudy (shamim):

Cost=0.08*3.14*1.5^2*7=?

OpenStudy (anonymous):

i dont no? sorry @shamim

OpenStudy (shamim):

Ok

OpenStudy (shamim):

Do u know cross sectional area of a cylinder

OpenStudy (shamim):

Do u know volume of cylinder

OpenStudy (anonymous):

i think its 3.26

OpenStudy (anonymous):

@shamim

OpenStudy (anonymous):

i have to take the test right now

OpenStudy (anonymous):

soooo am i right? @shamim

OpenStudy (shamim):

Right!!!

OpenStudy (anonymous):

can u help e with 1 more? @shamim

OpenStudy (shamim):

U should give me medal!

OpenStudy (shamim):

U can ask another as question in a new post

OpenStudy (anonymous):

ok